How much charge flows from each terminal of a 12.0-V battery when it is connected to a 4.10 μF capacitor?
Answer #1

if the capacitor is not initially charged then it wiil receive during its charging an electric charge Q given by
Q=V*C

Q=12*(4.10*10^-6) =49.2micro coulomb

However,

When charge reaches the value Q=49.2 micro culomco

​​​​ the electric field between the capacitor plates cancels the effect of the electric field generated by the battery, resulting in no further movement of charge.

Therefore, a capacitor when connected across a battery conducts for a while and after that behaves as an open circuit.